A biologist has a 30% solution and a 5% solution of the same plant nutrient. How many cubic centimeters (cc) of each solution should be mixed to obtain a 25 cc of a 25% solution? (Round your answers to 2 decimal places if needed)
Answer by nerdybill(7384)

A biologist has a 30% solution and a 5% solution of the same plant nutrient. How many cubic centimeters (cc) of each solution should be mixed to obtain a 25 cc of a 25% solution? (Round your answers to 2 decimal places if needed)
.
Let x = amount (cc) of 30% solution
then
25-x = amount (cc) of 5% solution
.
.30x + .05(25-x) = .25(25)
.30x + 1.25-.05x = 6.25
.25x + 1.25 = 6.25
.25x = 5
x = 20 cc (of 30% solution)
.
amount of 5% solution:
25-x = 25-20 = 5 cc